OpenStudy (anonymous):

A box contains 35 coins in quarters and nickels worth $3.15 How many of each coin are there (system of liner eqations problem)

OpenStudy (anonymous):

Q + N = 35 25Q + 5N = 315.

OpenStudy (anonymous):

Simplyf the second equation, solve the first for one of its unknowns and substitute into the simplified second equation, and you're well on your way to the solution.

OpenStudy (anonymous):

careful with the decimals! \[25q+5(35-q)=315\] \[

OpenStudy (anonymous):

\[20x+175=315\] \[20x=140\] \[x=7\]